The Jones family drove to Disneyland at 50 miles per hour and returned on the same route at 40mph. Find the distance to Disneyland if the total driving time was 7.2 hours.

Found 2 solutions by Nate, tom_from_nam:
Answer by Nate(3500)

distance = 50*t
distance = 40*(7.2 - t)
50t = 40(7.2 - t)
50t = 288 - 40t
90t = 288
t = 3.2
distance = 50*t or 160 miles

Answer by tom_from_nam(1)

First, we know the formula d = v x t, thus t = d/v.
The driving time from the Jones' house to Disneyland will equal d / 50 with d as the distance between their house and Disneyland.
Similarly, the driving time from Disneyland back to their house is d / 40.
According to the problem, we also know that the total time for driving to Disneyland and back is 7.2 hours.
Therefore, we can set up an equation with d being the unknown as follows:
(d/50) + (d/40) = 7.2
Multiplying both sides by 200, we get: 4d + 5d = 1440
Simplifying: 9d = 1440
Dividing both sides by 9, we get: d = 160
So the distance to Disneyland is 160 miles.